f(x) = (4x+6)
f (2) and f(8)

Answer by checkley77(12844)

f(x)=(4x+6)
f(2) and f(8)
f(2)=4*2+6
f(2)=8+6=14 ans.
f(8)=4*8+6
f(8)=32+6=38 ans.